# Policy Optimization with Smooth Guidance Learned from State-Only Demonstrations

### v2

@@ 7.2.2 Evaluation of Ground-Truth Reward Learning
A deeper look into the experimental results is provided to further verify the POSG’s effectiveness in learning near-optimal behaviors. In this experiment, for comparison, we trained the PPO algorithm with ground-truth environmental rewards that are the default setting of OpenAI Gym openaigym . We then calculated the average default reward of each step. Meanwhile, we deployed the experiment with POSG and other baseline methods in Fig. 4b in the sparse reward version of the same task. The default OpenAI Gym’s reward function was used to measure the quality of actions learned with these approaches, and then the average ground-truth reward for each step was computed. We compared the learning curves of the average ground-truth rewards for different algorithms. The experimental results are shown in Fig. 4c.
According to Fig. 4c, the POSG average ground-truth reward dramatically increases at the beginning of learning, similar to PPO’s learning trend in the default OpenAI Gym reward setting. At the same time, its value achieves an incremental improvement in the process of policy optimization. Moreover, the average ground-truth reward of POSG gradually approximates that of PPO in the default reward setting at the end of training. This result indicates that the smooth guidance reward can help the agent learn near-optimal behaviors that are rewarded a high score by the default OpenAI Gym reward function. For a more intuitive comparison, Fig. 5 shows an example of two policies learned using POSG with sparse rewards and PPO with dense default rewards. Interestingly, the POSG agent has learned gaits similar to those of PPO. This fact further demonstrates the effectiveness of POSG in learning a near-optimal policy in tasks with sparse rewards.
@@
Fig. 5: (a) Agent trained by POSG; (b) Agent trained by PPO with the ground-truth reward function.
7.3 Results in the Ant Maze
@@
Fig. 6: (a) Learning curves of success rate on the Ant Maze task; (b) Learning curves of average ground truth reward.
7.4 Ablation Analyses
@@
Fig. 7: (a) Learning curves of average return; (b) Learning curves of average ground truth reward.
7.4.1 Demonstrations with different amounts
@@
Fig. 8: (a) Learning curves of average return; (b) Learning curves of average ground truth reward.
7.4.2 Demonstrations with different qualities
@@ 8 Conclusion
Appendix A POSG Algorithm Training Process in Discrete Spaces
0:  learning rate $\alpha$, offline sparse-reward demonstration dataset $\mathcal{P}$, policy update frequency $K$
1:  Initialize policy weights $\theta$ of each agent.
4:  for $\text{episode}\in\{0,\dots,T\}$ do
6:     $R\leftarrow 0$    // Accumulate rewards for current episode
7:     for each step in $\{1,\dots,T\}$ do
8:        Choose $\boldsymbol{a}$ from $\boldsymbol{s}$ using $\pi_{\theta}$
11:     end for
14:     for each $((\boldsymbol{s},\boldsymbol{a}))$ in $\tau_{e}$ do
16:     end for
17:     if $\text{episode}\,\%\,K==0$ then
18:        Compute $A_{e}$ using environmental rewards
19:        Compute $A_{i}$ using guidance rewards
22:     end if
23:  end for
Algorithm 1 POSG in Discrete Spaces
Appendix B POSG Algorithm Training Process in High-Dimensional Continuous Spaces
0:  learning rate $\alpha$, offline sparse-reward demonstration dataset $\mathcal{P}$, policy update frequency $K$
1:  Initialize policy weights $\theta$ of each agent.
4:  for $\text{episode}\in\{0,\dots,T\}$ do
6:     $R\leftarrow 0$    // Accumulate rewards for current episode
7:     for each step in $\{1,\dots,T\}$ do
8:        Choose $\boldsymbol{a}$ from $\boldsymbol{s}$ using $\pi_{\theta}$
11:     end for
14:     if $\text{episode}\,\%\,K==0$ then
15:        Compute $A_{e}$ using environmental rewards
16:        Compute $A_{i}$ using guidance rewards
19:     end if
20:  end for
Algorithm 2 POSG in High-dimensional Continuous Spaces
Appendix C Proof of Lemma 1
@@
Appendix D Proof of Theorem 1 and Corollary 1
The following lemmas are proved in achiam2017constrained , and we have excerpted them here. The detailed proof process can be found in the appendix of achiam2017constrained .
@@ Appendix D Proof of Theorem 1 and Corollary 1
Acknowledgment
Conflict of interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper
Authors’ Contributions
Guojian Wang: Conceptualization, Methodology, Software, Formal analysis, Writing–Original Draft. Faguo Wu: Writing–Review and Editing, Supervision. Xiao Zhang: Validation, Supervision, Funding acquisition. Tianyuan Chen: Formal analysis, Software, Visualization. Zhiming Zheng: Validation, Supervision, Funding acquisition.
Availability of data and materials
The datasets generated during and/or analysed during the current study are available from the corresponding author on reasonable request.

### v3

@@ VII-B2 Evaluation of Ground-Truth Reward Learning
A deeper look into the experimental results is provided to further verify the POSG’s effectiveness in learning near-optimal behaviors. In this experiment, for comparison, we trained the PPO algorithm with ground-truth environmental rewards that are the default setting of OpenAI Gym [66]. We then calculated the average default reward of each step. Meanwhile, we deployed the experiment with POSG and other baseline methods in Fig. 5b in the sparse reward version of the same task. The default OpenAI Gym’s reward function was used to measure the quality of actions learned with these approaches, and then the average ground-truth reward for each step was computed. We compared the learning curves of the average ground-truth rewards for different algorithms. The experimental results are shown in Fig. 6.
@@
Figure 7: (a) Agent trained by POSG; (b) Agent trained by PPO with the ground-truth reward function.
According to Fig. 6, the POSG average ground-truth reward dramatically increases at the beginning of learning, similar to PPO’s learning trend in the default OpenAI Gym reward setting. At the same time, its value achieves an incremental improvement in the process of policy optimization. Moreover, the average ground-truth reward of POSG gradually approximates that of PPO in the default reward setting at the end of training. This result indicates that the smooth guidance reward can help the agent learn near-optimal behaviors that are rewarded a high score by the default OpenAI Gym reward function. For a more intuitive comparison, Fig. 7 shows an example of two policies learned using POSG with sparse rewards and PPO with dense default rewards. Interestingly, the POSG agent has learned gaits similar to those of PPO. This fact further demonstrates the effectiveness of POSG in learning a near-optimal policy in tasks with sparse rewards.
Figure 8: (a) Learning curves of success rate on the Ant Maze task; (b) Learning curves of average ground truth reward.
VII-C Results in the Ant Maze
@@
Figure 9: (a) Learning curves of average return; (b) Learning curves of average ground truth reward.
VII-D Ablation Analyses
@@
Figure 10: (a) Learning curves of average return; (b) Learning curves of average ground truth reward.
VII-D1 Demonstrations with different amounts
@@
VII-D2 Demonstrations with different qualities
@@ VIII Conclusion
Acknowledgments
Appendix A Proof of Lemma 1
@@
Appendix B Proof of Theorem 1 and Corollary 1
The following lemmas are proved in [62], and we have excerpted them here. The detailed proof process can be found in the appendix of [62].
@@ Appendix B Proof of Theorem 1 and Corollary 1
Appendix C POSG Algorithm Training Process in Discrete Spaces
Algorithm 1 POSG in Discrete Spaces
0:  learning rate $\alpha$, offline state-only demonstration dataset $\mathcal{P}$, policy update frequency $K$
1:  Initialize policy parameters $\theta$
4:  for $\text{episode}\in\{0,\dots,T\}$ do
6:     $R\leftarrow 0$    // Accumulate rewards for current episode
7:     for each step in $\{1,\dots,T\}$ do
8:        Choose $\boldsymbol{a}$ from $\boldsymbol{s}$ using $\pi_{\theta}$
11:     end for
14:     for each $((\boldsymbol{s},\boldsymbol{a}))$ in $\tau_{e}$ do
16:     end for
17:     if $\text{episode}\,\%\,K==0$ then
18:        Compute $A_{e}$ using environmental rewards
19:        Compute $A_{i}$ using guidance rewards
22:     end if
23:  end for
Appendix D POSG Algorithm Training Process in High-Dimensional Continuous Spaces
Algorithm 2 POSG in High-dimensional Continuous Spaces
0:  learning rate $\alpha$, offline state-only demonstration dataset $\mathcal{P}$, policy update frequency $K$
1:  Initialize policy parameters $\theta$
4:  for $\text{episode}\in\{0,\dots,T\}$ do
6:     $R\leftarrow 0$    // Accumulate rewards for current episode
7:     for each step in $\{1,\dots,T\}$ do
8:        Choose $\boldsymbol{a}$ from $\boldsymbol{s}$ using $\pi_{\theta}$
11:     end for
14:     if $\text{episode}\,\%\,K==0$ then
15:        Compute $A_{e}$ using environmental rewards
16:        Compute $A_{i}$ using guidance rewards
19:     end if
20:  end for